# 3Doodle: Compact Abstraction of Objects with 3D Strokes

### v1

@@ 4 Experiments
The inputs to 3Doodle are multi-view images and the corresponding camera poses. Our inputs are identical to recent neural scene reconstruction methods, and we evaluate the performance on the synthetic dataset from NeRF [29] and InvRender [51]. We additionally collect diverse shapes of synthetic objects from open-source 3D models. The input images are posed randomly on the upper hemisphere centered at the object and rendered with Blender’s Cycles path tracer [7]. We include samples from our dataset in Appendix B.
@@
Two of the baselines create 2D sketches from a single image. Kampelmühler and Pinz [18] train a class-specific generative network from sketch-image pairs, and works only on specific object categories. We therefore compare them only in balloon, car, chair, and teddy bear scenes. CLIPasso [40] spawns a composition of sparse 2D strokes from image input. We provide a single rendering of our 3D dataset for these baselines.
Three other baselines use 3D representations to draw sketch images with additional inputs. Among them, we categorize Artistic Radiance Fields (ARF) [47] and suggestive contours [9] to ones using detailed geometry to create dense lines. ARF first trains NeRF and converts the volume into the style of a reference image, to which we feed the result of our sketch. Suggestive contours extract geometric features from 3D mesh. Given the input images only, we reconstruct the mesh from state-of-the-art neural SDF reconstruction [41]. Lastly, Neural Edge Fields (NEF) [42] output sparse 3D primitives similar to our approach. It produces geometric feature curves in 3D given multi-view 2D edge maps, estimated from PiDiNet [39].
4.1 Quantitative Evaluation
Table 1: Sketch recognition accuracy. We mark the best and second-best results in bold and underlined numbers.
@@ 4.1 Quantitative Evaluation
Figure 4: Bar plots of the perceptual study results.
@@
The answers from 72 participants are summarized in Figure 4. 3Doodle receives significantly higher score for being realistic sketches compared to the baselines using dense representations (ARF and suggestive contours). The results also indicate that the participants appreciate the effectiveness of our sparse representation. Dense representations, on the other hand, can preserve the fine 3D geometry and maintain highly expressive and consistent structure, as expected. 3Doodle achieves a comparable score in expressiveness despite its compactness, which is a great improvement from other sparse representations (NEF and CLIPasso). CLIPasso marks the lowest score in multi-view consistency, while 3Doodle maintains inherent view consistency that comes from the 3D representation.
@@
Figure 5: Qualitative results. We show the target objects in the leftmost column and multi-view rendered results of our 3D strokes in the right three columns.
4.2 Qualitative Evaluation
@@
We demonstrate the qualitative results of our sketches in Figures 1 and 5. 3Doodle successfully expresses the essential structures of various types of objects with a highly compact representation. Our view-independent sketches $\mathcal{S}^{\text{3D}}_{\text{ind}}$ capture both geometric features (e.g. fine structures of sails in the boat scene) and semantic features (e.g. a wave pattern in the boat scene). Also, our view-dependent components $\mathcal{S}^{\text{3D}}_{\text{dep}}$ successfully represent the smooth bounding surfaces that envelop the objects. While $\mathcal{S}^{\text{3D}}_{\text{dep}}$ expresses the outline of objects (e.g. shapes of teddy bear and snowman), $\mathcal{S}^{\text{3D}}_{\text{ind}}$ expresses the fine details from texture and sharp geometry (e.g. arms of snowman and face of teddy bear) to draw whole sketches together.
@@
We provide a visual comparison between baselines with 3D representation in Figure 12. The sketches generated by suggestive contours are close to shading rather than human free-hand line drawing since the surfaces of reconstructed 3D meshes are not perfect as human assets. Also, ARF, which converts the color of pre-trained plenoxels [12], draws random strokes on the surface of objects (e.g. the random strokes on the plate of the hotdog scene and the surface of the teddy bear). In contrast, NEF and 3Doodle produce more human-like drawings since both methods represent sketches as a set of geometric primitives. However, NEF fails to represent the objects which contain smooth surfaces (teddy bear and snowman) since their edge fields only extract the view-independent parametric curves. In contrast to NEF, 3Doodle successfully represents the sketch of the smooth surfaces with $\mathcal{S}^{\text{3D}}_{\text{dep}}$.
@@
We further visually compare our approach with image sketch generation algorithms in Figure 13. The approach by Kampelmühler and Pinz are not designed to generate sketches for unseen categories (e.g. lego scene) since they need training from sketch-image paired datasets. Also, the sketches generated for the seen categories are not faithful to the target objects. The optimization-base methods CLIPasso and 3Doodle successfully produce sketches that abstract a wide variety of objects with a sparse set of strokes. However, CLIPasso fails to maintain consistency across the different viewpoints. (e.g. the vertical lines of the air balloon, lines on the chair seat, and the base of the Lego) On the contrary, 3Doodle renders view-consistent 2D sketches due to the 3D representation.
@@
Figure 6: 3Doodle robustly generates sketch for a few multi-view inputs.
Furthermore, our compact 3D primitives have a small number of parameters, providing an additional advantage of robustly generating sketches even from a small number of input images. In Figure 6, we demonstrate the results of our sketch generated from a smaller number of input views. We observe that we can obtain a reasonable sketch from only 15 views which is much smaller than the default 100 views of the NeRF training dataset.
@@
Figure 7: Sketch stylization. We demonstrate various styles of our generated sketch by rendering $\mathcal{S}^{\text{3D}}$ as vector graphics format and applying various brush styles from Adobe Illustrator.
@@
Also, one can alter the style of our sketch by applying different brushes to the vector strokes as illustrated in Figure 7.
4.3 Ablation Studies
@@
Figure 8: Effects of sketch components. The view-independent sketches cannot solely represent the objects that have smooth surfaces.
@@
In this section, we analyze the effects of important components of 3Doodle. First, we study about the necessity of the view-dependent component $\mathcal{S}^{\text{3D}}_{\text{dep}}$, which is our novel addition compared to previous works employing line-only representations. The volumetric encapsulation is critical especially for sphere objects, as in Figure 8. When we only use 3D line strokes ($\mathcal{S}^{\text{3D}}_{\text{ind}}$) the 3D primitives clearly fail to express the contours of smooth surfaces. NEF also uses view-independent 3D strokes, and exhibits similar characteristics when processing round objects with view-dependent contours, such as teddy bear and snowman in Figure 12.
Figure 9: Effects of loss function.
Table 2: Quantitative results on ablated loss terms
@@
Then we evaluate the individual terms in our loss term. As shown in Figure 9, both LPIPS and CLIP losses are crucial to effectively capture perceptual characteristics with sparse primitives. Without LPIPS, the resulting sketch deviates from the given geometric structure. On the other hand, removing CLIP loss retains the overall layout of the input object, but misses essential parts to be a ‘chair’. We measure our quantitative metrics of the ablated versions with the chair object in Table 2. Interestingly, removing CLIP loss does not necessarily achieve better LPIPS loss, or vice versa. Our full loss function can find the better optimal point for both terms. We also test a devised loss function which optimizes line strokes to match edge maps extracted from images [39]. All evaluation metrics are inferior, and the resulting sketch shows remaining short, un-optimized strokes. We can infer that the loss directly employing detailed edge-based images suffers from a complex convergence basin with local minima. Our full loss term efficiently finds an abstract sketch of an object.
Figure 10: Effects of initialization step.
Next, we demonstrate the effectiveness of initializing our 3D primitives from SfM points instead of random initialization. In Figure 10, the top row shows the results using view-dependent sketches $\mathcal{S}^{\text{3D}}_{\text{dep}}$, and the bottom row contains the view-independent sketches $\mathcal{S}^{\text{3D}}_{\text{ind}}$. Our optimization objective is highly non-convex, and the sketch parameters for both $\mathcal{S}^{\text{3D}}_{\text{dep}}$ and $\mathcal{S}^{\text{3D}}_{\text{ind}}$ fail to fully cover the shape of target objects without a proper initialization strategy. In particular, we observe that several strokes are condensed to partially cover high-contrast regions.
Figure 11: Different abstraction levels by modifying the number of sketch components.
We also study the effects of the number of strokes in 3Doodle. In Figure 11, we display the different abstraction levels of sketches by varying the number of Bézier curves. Overall, 3Doodle can generate high-quality sketches with strokes ranging from 16 to 48 strokes. As the number of strokes increases, 3Doodle effectively captures finer details, whereas a higher level of abstraction is achieved in expressing the object with fewer strokes. However, when the number of Bézier curves is too small compared to the complexity of the input object, 3Doodle may fail to represent the geometric details of the target object. (e.g., the right-side armrest of the chair is erased using only 16 strokes in Figure 11.)
5 Conclusion
@@
Figure 12: Qualitative comparison with baseline methods which have 3D representations.
Figure 13: Qualitative comparison with image sketch generation methods.
@@ Appendix A Projection of 3D Bézier curve
A.1 Orthogonal Projection
We provide proof of Theorem 1 in this subsection. See 1
@@ A.1 Orthogonal Projection
$\tilde{B}^{\text{2D}}(t)=\sum_{j=0}^{3}b_{j}(t)\begin{pmatrix}p_{x}^{j}\\ p_{y}^{j}\\ 0\end{pmatrix},$ (12)
@@
A.2 Perspective Projection
In this subsection, we provide proof of Theorem 2, which describes the property of perspective projection of 3D cubic Bézier curve. Then, we explain when our approximation is reasonable.
@@ A.2 Perspective Projection
$\tilde{B}^{\text{2D}}(t)$ $=\begin{pmatrix}B^{\text{3D}}_{x}(t)\dfrac{f}{B^{\text{3D}}_{z}(t)}\\ B^{\text{3D}}_{y}(t)\dfrac{f}{B^{\text{3D}}_{z}(t)}\end{pmatrix}$ (15)
$=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)fp_{x}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)fp_{y}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\end{pmatrix}$ (18)
$=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)\left(\frac{f}{p_{z}^{j}}p_{x}^{j}\right)p_{z}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)\left(\frac{f}{p_{z}^{j}}p_{y}^{j}\right)p_{z}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\end{pmatrix}$ (21)
$=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)q_{x}^{j}\omega_{j}}{\sum_{j=0}^{3}b_{j}(t)\omega_{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)q_{y}^{j}\omega_{j}}{\sum_{j=0}^{3}b_{j}(t)\omega_{j}}\end{pmatrix},$ (24)
@@
Following Theorem 2, our Bézier curve approximation of the rational Bézier curve is valid when the weights assigned to four control points are nearly the same. Namely, our approximation is reasonable when the viewpoints are located sufficiently far from the interested objects.

### v2

@@ 4. Experiments
The inputs to 3Doodle are multi-view images and the corresponding camera poses. Our inputs are identical to recent neural scene reconstruction methods, and we evaluate the performance on the synthetic dataset from NeRF (Mildenhall et al., 2021) and InvRender (Zhang et al., 2022b). We additionally collect diverse shapes of synthetic objects from open-source 3D models. We render input images with Blender’s Cycles path tracer, whose views are posed randomly on the upper hemisphere, centered at the object. We include samples from our dataset in Appendix B. We also test our method on the real-world scenes from the CO3D dataset (Reizenstein et al., 2021).
@@
Two of the baselines create 2D sketches from a single image. Kampelmühler and Pinz (2020) train a class-specific generative network from sketch-image pairs, and works only on specific object categories. We therefore compare them only in balloon, car, chair, and teddy bear scenes. CLIPasso (Vinker et al., 2022) spawns a composition of sparse 2D strokes from image input. We provide a single rendering of our 3D dataset for these baselines.
Three other baselines use 3D representations to draw sketch images with additional inputs. Among them, we categorize Artistic Radiance Fields (ARF) (Zhang et al., 2022a) and suggestive contours (DeCarlo et al., 2003) to ones using detailed geometry to create dense lines. ARF first trains NeRF and converts it into the style of a reference image, to which we feed the result of our sketch. Suggestive contours extract geometric features from 3D mesh. Given the input images only, we reconstruct the mesh from state-of-the-art neural SDF reconstruction (Wu et al., 2023). Lastly, Neural Edge Fields (NEF) (Ye et al., 2023) output sparse 3D primitives similar to our approach. It produces geometric feature curves in 3D given multi-view 2D edge maps, estimated from PiDiNet (Su et al., 2021). However, NEF fails to reconstruct meaningful edge fields in our toycar and toyhorse scenes so we did not include the results.
4.1. Quantitative Evaluation
Table 1. Sketch recognition accuracy. We mark the best and second-best results in bold and underlined numbers.
@@ 4.1. Quantitative Evaluation
Figure 4. Bar plots of the perceptual study results.
@@
The answers from 72 participants are summarized in Fig. 4. 3Doodle receives significantly higher score for being realistic sketches compared to the baselines using dense representations (ARF and suggestive contours). The results also indicate that the participants appreciate the effectiveness of our sparse representation. Dense representations, on the other hand, can preserve the fine 3D geometry and maintain highly expressive and consistent structure, as expected. 3Doodle achieves a comparable score in expressiveness despite its compactness, which is a great improvement from other sparse representations (NEF and CLIPasso). CLIPasso marks the lowest score in multi-view consistency, while 3Doodle maintains inherent view consistency that comes from the 3D representation.
@@
Figure 5. Qualitative results. We show the target objects in the leftmost column and multi-view rendered results of our 3D strokes in the right three columns.
4.2. Qualitative Evaluation
@@
We demonstrate the qualitative results of our sketches in Figs. 1 and 5. 3Doodle successfully expresses the essential structures of various types of objects with a highly compact representation. Our view-independent sketches $\mathcal{S}^{\text{3D}}_{\text{ind}}$ capture both geometric features (e.g. fine structures of sails in the boat scene) and semantic features (e.g. a wave pattern in the boat scene). Also, our view-dependent components $\mathcal{S}^{\text{3D}}_{\text{dep}}$ successfully represent the smooth bounding surfaces that envelop the objects. While $\mathcal{S}^{\text{3D}}_{\text{dep}}$ expresses the outline of objects (e.g. shapes of teddy bear and snowman), $\mathcal{S}^{\text{3D}}_{\text{ind}}$ expresses the fine details from texture and sharp geometry (e.g. arms of snowman and face of teddy bear) to draw whole sketches together.
@@
We provide a visual comparison between baselines with 3D representation in Fig. 12. The sketches generated by suggestive contours are close to shading rather than human free-hand line drawing since the surfaces of reconstructed 3D meshes are not perfect as human assets. Also, ARF, which converts the color of pre-trained plenoxels (Fridovich-Keil et al., 2022), draws random strokes on the surface of objects (e.g. the random strokes on the plate of the hotdog scene and the surface of the teddy bear). In contrast, NEF and 3Doodle produce more human-like drawings since both methods represent sketches as a set of geometric primitives. However, NEF fails to represent the objects which contain smooth surfaces (teddy bear and snowman) since their edge fields only extract the view-independent parametric curves. In contrast to NEF, 3Doodle successfully represents the sketch of the smooth surfaces with $\mathcal{S}^{\text{3D}}_{\text{dep}}$.
@@
We further visually compare our approach with image sketch generation algorithms in Fig. 13. The approach by Kampelmühler and Pinz are not designed to generate sketches for unseen categories (e.g. lego scene) since they need training from sketch-image paired datasets. Also, the sketches generated for the seen categories are not faithful to the target objects. The optimization-base methods CLIPasso and 3Doodle successfully produce sketches that abstract a wide variety of objects with a sparse set of strokes. However, CLIPasso fails to maintain consistency across the different viewpoints. (e.g. the vertical lines of the air balloon, lines on the chair seat, and the base of the Lego) On the contrary, 3Doodle renders view-consistent 2D sketches due to the 3D representation.
@@
Figure 6. 3Doodle robustly generates sketch for a few multi-view inputs.
@@
Figure 7. Sketch stylization. We demonstrate various styles of our generated sketch by rendering $\mathcal{S}^{\text{3D}}$ as vector graphics format and applying various brush styles from Adobe Illustrator.
Furthermore, our 3D sketch have a small number of parameters, providing an additional advantage of robustly generating sketches even from a small number of input images. In Fig. 6, we demonstrate the results of our sketch generated from a smaller number of input views. We also uniformly sample the viewpoints on the upper hemisphere. We observe that we can obtain a reasonable sketch from only 15 views which is much smaller than the default 100 views of the NeRF training dataset. Our approach can also generate sketches from real-world captures in CO3D (Reizenstein et al., 2021) without further adjustments as shown in the last two columns in Fig. 5.
@@
Also, one can alter the style of our sketch by applying different brushes to the vector strokes as illustrated in Fig. 7.
4.3. Ablation Studies
@@
Figure 8. Effects of sketch components. The view-independent sketches cannot solely represent the objects that have smooth surfaces.
@@
In this section, we analyze the effects of important components of 3Doodle. First, we study about the necessity of the view-dependent component $\mathcal{S}^{\text{3D}}_{\text{dep}}$, which is our novel addition compared to previous works employing line-only representations. The volumetric encapsulation is critical especially for sphere objects, as in Fig. 8. When we only use 3D line strokes ($\mathcal{S}^{\text{3D}}_{\text{ind}}$) the 3D primitives clearly fail to express the contours of smooth surfaces. NEF also uses view-independent 3D strokes, and exhibits similar characteristics when processing round objects with view-dependent contours, such as teddy bear and snowman in Fig. 12.
Figure 9. Effects of loss function.
Table 2. Quantitative results on ablated loss terms
@@
Then we evaluate the individual terms in our loss term. As shown in Fig. 9, both LPIPS and CLIP losses are crucial to effectively capture perceptual characteristics with sparse primitives. Without LPIPS, the resulting sketch deviates from the given geometric structure. On the other hand, removing CLIP loss retains the overall layout of the input object, but misses essential parts to be a ‘chair’. We measure our quantitative metrics of the ablated versions with the chair object in Tab. 2. Interestingly, removing CLIP loss does not necessarily achieve better LPIPS loss, and vice versa. Our full loss function can find the better optimal point. We also test a devised loss function which optimizes line strokes to match edge maps extracted from images (Su et al., 2021). All evaluation metrics are inferior, and the resulting sketch shows remaining short and un-optimized strokes. We can infer that the loss directly employing detailed edge-based images suffers from a complex convergence basin with local minima. Our full loss term efficiently finds an abstract sketch of an object.
Figure 10. Effects of initialization step.
Next, we demonstrate the effectiveness of initializing our 3D primitives from SfM points in Fig. 10. As a comparison, we show results where the control points and the superquadric centers are initialized randomly or using the output of line-based SfM (Liu et al., 2023). Our approach is stable and produces a similar quality of results in most scenes. However, our non-convex optimization relies on image observations and may condense 3D strokes in high-contrast regions or miss details without much image evidence, such as a white arm of a chair against a white background. In such challenging scenarios, SfM point clouds provide a proper initialization to fully cover the shape of the target objects with a minimal number of primitives, such as superquadrics in the teddy bear. Initialization from line-based SfM can partially mitigate the problem, especially for structures that can be approximated by straight lines, and sometimes results in faster convergence if we only use Bézier curves $\mathcal{S}^{\text{3D}}_{\text{ind}}$. However, it struggles to reconstruct curvy boundaries in a teddy bear or a rocking pony. The details on the initialization pipeline can be found in Appendix.
Figure 11. The effects of using different numbers of sketch components $N_{\text{ind}}$. With more strokes, 3Doodle captures finer details. Fewer strokes lead to more abstract expression.
We also study the effects of the number of strokes in 3Doodle. In Fig. 11, we display the different abstraction levels of sketches resulted by varying the number of Bézier curves. Overall, 3Doodle can generate high-quality sketches with strokes ranging from 16 to 48 strokes. As the number of strokes increases, 3Doodle effectively captures finer details, whereas a higher level of abstraction is achieved in expressing the object with fewer strokes. However, when the number of Bézier curves is too small compared to the complexity of the input object, 3Doodle may fail to represent the geometric details of the target object (e.g., the right-side armrest of the chair is erased using only 16 strokes in Fig. 11).
5. Conclusion and Future Works
@@
3Doodle has many interesting potential applications. As an automatic way to generate multiple sketch images of the 3D shape, 3Doodle can build a dataset of 3D sketch-images or 3D sketch-shape pairs, which is more scalable than having humans draw the sketches (Luo et al., 2021). The fully differentiable pipeline can be easily extended to creative content creations beyond multi-view renderings, such as text-to-3D sketch generation or 3D sketch video generation. Furthermore, the compact structural proxy may assist other practical applications that require 3D geometric information. Future works include using the approximate 3D shape to guide recognition, localization (Liu et al., 2023; Fabbri and Kimia, 2010), few-shot NeRF (Deng et al., 2022; Guangcong et al., 2023), or other 3D reconstruction tasks, where geometric or depth priors are known to accelerate and stabilize the outcome.
@@
Acknowledgements.
This work was supported by the NRF grants (No.RS-2023-00208197 (50%) and No.2023R1A1C200781211(30%)), IITP grant funded by the Korea government(MSIT) (No.2021-0-01343, AI Graduate School Program(Seoul National University)(10%)) funded by the Korea government(MSIT), INMC, and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024(10%). Young Min Kim is corresponding author.
@@ Appendix A Projection of 3D Bézier curve
A.1. Orthogonal Projection
We provide proof of 1 in this subsection. See 1
@@ A.1. Orthogonal Projection
(12) $\tilde{B}^{\text{2D}}(t)=\sum_{j=0}^{3}b_{j}(t)\begin{pmatrix}p_{x}^{j}\\ p_{y}^{j}\\ 0\end{pmatrix},$
@@
A.2. Perspective Projection
In this subsection, we provide proof of 2, which describes the property of perspective projection of 3D cubic Bézier curve. Then, we explain when our approximation is reasonable.
@@ A.2. Perspective Projection
(13) $\tilde{B}^{\text{2D}}(t)$ $=\begin{pmatrix}B^{\text{3D}}_{x}(t)\dfrac{f}{B^{\text{3D}}_{z}(t)}\\ B^{\text{3D}}_{y}(t)\dfrac{f}{B^{\text{3D}}_{z}(t)}\end{pmatrix}$
(14) $=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)fp_{x}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)fp_{y}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\end{pmatrix}$
(15) $=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)\left(\frac{f}{p_{z}^{j}}p_{x}^{j}\right)p_{z}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)\left(\frac{f}{p_{z}^{j}}p_{y}^{j}\right)p_{z}^{j}}{\sum_{j=0}^{3}b_{j}(t)p_{z}^{j}}\end{pmatrix}$
(16) $=\begin{pmatrix}\dfrac{\sum_{j=0}^{3}b_{j}(t)q_{x}^{j}\omega_{j}}{\sum_{j=0}^{3}b_{j}(t)\omega_{j}}\\ \dfrac{\sum_{j=0}^{3}b_{j}(t)q_{y}^{j}\omega_{j}}{\sum_{j=0}^{3}b_{j}(t)\omega_{j}}\end{pmatrix},$
@@
Following 2, our Bézier curve approximation of the rational Bézier curve is valid when the weights assigned to four control points are nearly the same. Namely, our approximation is reasonable when the viewpoints are located sufficiently far from the interested objects.